OpenStudy (anonymous):

-19+(-6)=? How do I solve this?

OpenStudy (anonymous):

(-25) sure....^_^

OpenStudy (anonymous):

-19-6 =-25 :)

OpenStudy (anonymous):

when you add a negative number, you are in effect subtracting it. so -19+(-6) becomes -19-6

OpenStudy (anonymous):

:D Ok....so if I had another problem like -12+(-9), it would be -12-9=-21? :)

OpenStudy (anonymous):

You take away the addation and () and then when you solve the equation it = -25 (:

OpenStudy (anonymous):

yes jessica, very good.

OpenStudy (anonymous):

yes ,, you are correct.. (-) + (-) = -

OpenStudy (anonymous):

yes it would. with the smiley at the end of course.

OpenStudy (anonymous):

Lol

OpenStudy (anonymous):

haha man smileys are the best :)

OpenStudy (anonymous):

They Def are!! (:

OpenStudy (anonymous):

Ok, mahalo nui loa to all of you. ^___^ BTW Fiddle look----> (◣‿◢) It's my new face! :D

OpenStudy (anonymous):

It's so you !!! :)

OpenStudy (anonymous):

^__^ Thank you, Fiddle.